OpenStudy (anonymous):

Please Help Use The Function Below To Find F(4) F(x)=3^x A.27 B.12 C.7 D.81

OpenStudy (anonymous):

im not sure if im suppose to do 3x3x3?

OpenStudy (anonymous):

no to the power of 4, not 3

OpenStudy (isaidavila):

3^(4)=?

OpenStudy (anonymous):

so it will be 81 am i correct? @satellite73 @isaidavila

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 3^x\] \[\Large f({\color{red}{x}}) = 3^{{\color{red}{x}}}\] \[\Large f({\color{red}{4}}) = 3^{{\color{red}{4}}}\] \[\Large f({\color{black}{4}}) = 3*3*3*3\] \[\Large f({\color{black}{4}}) = 81\] You have the correct answer

OpenStudy (anonymous):

thank you all so much i appreciate it!